OpenStudy (anonymous):

Solve for V in V = s3, if s = 4.

OpenStudy (johnweldon1993):

\[V = s^3\] I'm assume that is what the question looks like?

OpenStudy (anonymous):

Yes.

OpenStudy (johnweldon1993):

So you have a value for 's'....that you can just plug into that equation meaning \[V = 4^3\] what is 4 to the third power? 4 * 4 * 4 = ?